# Improving Technical ”How-to” Query Accuracy with Automated Search Results Verification and Reranking

### v1

@@ 5.2. Further Analysis and Ablation Studies
Despite the promising performance shown in the zero-shot setting, we want to further explore the correctness based on our data setting and in which cases the system stumbled. Therefore, we conducted further analysis in this section to answer several critical questions.
@@
To answer this question, an analysis of both success and failure cases are conducted to measure the failure rate at each stage. After the instruction extraction stage, 398 pages (11.98%) have extracted instructions from the total 3321 pages, with 117 queries (about 70.05%) out of the total 167 queries. Among the 398 pages, 75 pages (18.8%) with 39 queries(33.33%) contain relevant instructions, while the remaining 323 pages with 108 queries are negative samples. At this stage, LLMs’ hallucination issue is a major factor: generative content unaligned with original page content induces noises, which decreases the percentage of true positive pages. We will further analyze how the extraction performance has been impacted by HTML grounding in question 2.
At the execution stage, the execution engine filters out 119(29.90%) pages from 68(58.12%) queries and only 279 pages from 91 queries are left. For those remaining pages, if the execution engine can finish more than one instruction or the instruction completion degree of the corresponding execution is larger than 0, the extracted instructions are considered as executable and are forwarded to the next stage. Among those dropped, 18(15.12%) pages are ones with no action taken by the execution engine and 101 (84.88%)pages have a score of 0 generated by GPT4-V. Considering the result of the former is promising, we focus on analysis of : clearly shown in table 6, about 25% of pages with no or partial instructions are set with = 0 and about 70% makes the execution engine struck at the first instruction (the engine can take actions on mobile screen but can’t move forward due to the missing necessary information). Another good indicator is that about 75% pages with instructions can support the execution engine to finish more than 75% instructions, which is aligned with human verification. However, top 25% pages with no or partial instructions are mistakenly marked as = 1, introducing significant noise in reranking. This highlights the necessity of drawing more robust, execution-derived signals from the execution in future work.
@@
At the reranking stage, we analyze the relationship between features and reranking scores for the best-performed models. Most successful cases are because models figure out strong signals among input features, such as and , and learn accurate non-linear weights to reflect on their interrelationship. However, current models tend to minimize negative signals even for those with abnormal values: for instance, a feature vector with several very low values, like 2.8e-05, will still be assigned a higher score just because of multiple features that contain higher values. We attribute this to insufficient negative examples of training and validation sets or limited features incorporated in reranking.
@@
We use page URLs collected in ”How-to” WeWeb dataset to fetch the corresponding web pages and apply instruction extraction on individual pages to compare whether the extracted content matches the ground truth steps. Accuracy is used to evaluate the instruction extraction performance, which measures whether the model appropriately extracts steps or outputs ”none” based on the given web page and search query. GPT4 extraction accuracy is compared with and without step grounding (Table 7). The results show IE without grounding (full or partially) is pages out of 169 web pages with relevant instructions. 42 pages with instructions were partially extracted. However, it also falsely extracted instructions from 953 pages that don’t contain relevant instructions. This happens because the model can hallucinate, extract the wrong set of instructions from a web page with multiple sets of instructions, or significantly paraphrase the instructions.
With grounding, many partially extracted instructions are corrected by adding missing steps on the same XPaths on the HTML DOM tree, and hallucinated instructions are removed. As a result, grounding greatly reduced the number of false positive (i.e. extracted instructions from pages without relevant instructions) from 953 to 303, meanwhile reducing the number of true positive (i.e. extracted instructions from pages with relevant instructions) from to .
@@
We expect LTR will work if we have training data with the same distribution as the test data. However, considering the huge difference between the pre-training and testing data, is it still valid to have a LTR component? We tried to rerank by a simple rule: promoting pages with (i.e. GPT4-V predicts the instructions are 100% completed) to the top without changing their related rankings. This rule has a , significantly worse than the proposed approach and not significantly better than the baseline. This suggests the LTR module is more useful than the simple rule.
Q4. How sensitive the model performance is for different LTR model settings?
We varied the TMLP model structure by adding Sigmoid, Tanh activation functions. We found the model performance changed a little. However, there is nothing statistically significant. We expect the small size of our test dataset and the big gap between training and testing data causes the results less sensitive to the perturbations on TMLP. The LTR models we trained are far from optimal, and we expect better training data could improve the performance.
@@ 5.3. Discussion and Future Work
While our study represents a novel approach for search ranking, we recognize that this is a preliminary study of the problem, and as a result comes with certain limitations.
@@
One of the primary limitations of our study is that it has been done only on the Android platform, and other platforms such as desktop and web browser are not covered and should be studied in the future. Although we cover a diverse set of domains, the coverage is not exhaustive. Our study also shows the exact performance (P@1 etc.) is domain-specific. This may affect the generalization ability of our findings and the reader should take this into account when interpreting the results. When applying similar techniques to other platforms or domains, the exact number (P@1 etc.) would be different.
@@
Although generative LLMs offer a rapid start, the extraction model is significantly hindered by hallucination issues, as noted by Ji et al. (Ji et al., 2023). Our current methodology, which focuses on post-processing with grounding, proves to be a weak regularizer. Moving forward, we would like to try generative LMs with multimodal document models such as XDoc (Chen et al., 2022), MarkupLM (Li et al., 2022b), and LayoutLMv3 (Huang et al., 2022b), aiming to enhance in-model grounding efficacy.
@@
A more tightly coupled integration of the UI context and the Action Prediction Model can further enhance the execution performance. Moreover, our current prompt-based method which used the actionable regions as input does not effectively represent semantic groups on the UI layer. For instance, an input box and its label username should be semantically grouped as one region for the execution module. Another next step is to enhance the previous work (Hu, 2020), (Liu et al., 2018) to improve the UI contextual information passed to the Action Prediction Model.
@@
The reranking model we used is very preliminary, using simple feature engineering and no real-world training data. With more resources, we should collect more real-world training data with strategic sampling. Instead of using manually engineered features, we could represent the verification information, especially the sequence of action screens and instructions, through supervised or unsupervised learning. With better reranking models, we would expect the effectiveness of the proposed approach to be more significant.
@@
Although verification agents can save a user time, it might accidentally take harmful actions as some online content could be misleading. In practice, we need to implement safeguards to prevent harmful actions from being taken by the execution agent. Ideally, a client environment that doesn’t impact the user’s real environment should be used. If not, we might let each user review and approve extracted instructions or at-risk actions before the automated execution. Another approach is to run server-side verification offline and display metadata about verification in the search results (e.g. ”verified for app1 v3 on android v11”), which allows personalized search results without real time verification on the client side.
@@ 6. Conclusion
We propose reranking top retrieved results for ”How-to” search queries by promoting candidate web pages with verified executable instructions, which adds an additional layer of usefulness to a typical search engine workflow. It ranks search results based on the actual completeness of the candidate instructions, besides traditional metrics such as textual relevance and authority scores. The experimental results demonstrate the proposed approach could further improve a very strong baseline search engine(i.e. Google).
This paper is a first step forward in improving the reliability and usefulness of search results of online help resources for ”How-to” queries about software tasks. There are several future directions as discussed before, like adapting the system 1 to support verification of other platforms (web, macOS, Windows, iPhone etc.), or collecting user feedback to improve the reranking models and also improving the software agent’s ability to interpret and execute a wider range of instructions.
@@ Acknowledgements.
For ”How-to” WeWeb dataset, we appreciate our research collaborators Sijia Zhong, Valentina Tang, Zoey Zhou, Jimmy Chen, and Chole Wong. Without their efforts, it is impossible to finish collecting execution data in a short time period. They gave valuable suggestions on MagicWand, making it more mature to support our research work. Thank Yaxuan Wang for her efforts to refine figures used in this paper, helping illustrate key ideas concisely and elegantly. Annabelle Miin, Krish Pai, Rithvik Chavali, Saket Pathak have involved this research from the beginning and finished work, like deciding on Android Apps, search queries, and guaranteeing data quality via two phases verification.References

### v3

@@ 5.2. Further Analysis and Ablation Studies
Despite the promising performance on the zero-shot setting, we want to further explore the correctness based on our data setting and in which cases the system stumbled. In this section, we conduct a further analysis to answer several critical questions.
@@
To answer it, success and failure cases have been analyzed to measure the failure rate at each stage. After the instruction extraction stage, 398 pages (11.98%) have extracted instructions from the total 3321 pages, with 117 queries (about 70.05%) out of the total 167 queries. Among the 398 pages, 75 pages (18.8%) with 39 queries(33.33%) contain relevant instructions, while the remaining 323 pages with 108 queries are negative samples. At this stage, LLMs’ hallucination issue is a major factor: generative content unaligned with original page content induces noises, which decreases the percentage of true positive pages. We will further analyze how HTML grounding has impacted extraction performance in Q2.
At the execution stage, the execution engine filters 119(29.90%) pages from 68(58.12%) queries, and only 279 pages from 91 queries are left. For the remaining pages, if the execution engine can complete more than one instruction or the instruction completion degree of the corresponding execution is greater than 0, the extracted instructions are considered executable and are advanced to the next stage. Among those dropped, 18(15.12%) pages have no action taken by the execution proxy, and 101 (84.88%)pages have a score of 0. Considering the result of the former is promising, we focus on an analysis based on : clearly shown in table 6, about 25% of pages with no or partial instructions are set with = 0 and about 70% makes the execution engine struck at the first instruction (the engine can’t move forward due to the missing necessary information). Another good indicator is that about 75% pages with instructions can support the execution engine to finish more than 75% instructions, which is aligned with human verification. However, top 25% pages with no or partial instructions are mistakenly marked as = 1, introducing significant noise to reranking. This highlights the need to incorporate more robust, execution-derived signals in future work.
@@
In the re-ranking stage, we analyze the relationship between features and reranking scores for the best-performed models. Most successful cases occur when models identify strong signals among input features, such as and , and accurately learn non-linear weights to reflect on their interrelationship. However, current models tend to minimize negative signals even those with abnormal values: for instance, a feature vector containing several very low values, like 2.8e-05, will still receive a higher score just if it includes multiple features with higher values. We attribute this to insufficient negative examples of training and validation sets or limited features incorporated in reranking.
@@
We use page URLs collected in the ”How-to” WeWeb dataset to fetch the corresponding web pages and apply instruction extraction on individual pages to compare whether the extracted content matches the ground-truth steps. Accuracy is used to evaluate the instruction extraction performance, which measures whether the model appropriately extracts steps or outputs ”none” under different cases. Extraction accuracy is listed with and without step grounding (Table 7). The results show that instruction extraction without grounding (full or partial) is out of 169 web pages with relevant instructions. 42 pages with instructions were partially extracted. However, it also falsely extracted instructions from 953 pages that don’t contain relevant instructions. This happens because the model can hallucinate, extract the wrong instructions from a web page with relevant and irrelevant instructions, or significantly paraphrase the instructions.
With grounding, many partially extracted instructions are corrected by adding missing steps on the same XPaths on the HTML DOM tree, removing hallucinated instructions. As a result, it greatly reduced the number of false positives (i.e., extracted instructions from pages without relevant instructions) from 953 to 303, meanwhile reducing the number of true positives (i.e., extracted instructions from pages with relevant instructions) from to .
@@
We expect LTR will work better if we have training data with a similar distribution as the test data. However, considering the huge difference between the pre-training and testing data, is it still valid to have an LTR component? To clarify this question, we tried to rerank using a simple rule: promoting pages with (i.e. GPT4-V predicts the instructions are 100% completed) to the top without changing their related rankings. This rule-based reranking reaches , much worse than our proposed LTR approaches while not significantly better than the baseline. This suggests LTR models are more useful than a simple rule based on one strong feature.
Q4. Which features contribute more to the overall performance?
During the reranking performance analysis, we found that significantly influenced the final ranking: LTR models tend to assign a higher ranking score when values are higher. This tendency is more obvious when and are aligned. We attribute it to the limited diversity of our synthetic training data, leading to ranking decisions being dominated by strong features. Another interesting observation is that the remaining features, which focus on instruction, UI, and action alignments, fail to address some common but tricky circumstances in agent execution, resulting in incorrect ranking decisions. For instance, agents may get trapped in ”execution cycles” where robots revert to a previous state due to ambiguous instructions. Despite this, the visual features and LTR models still tend to assign higher scores to these instructions. This underscores the necessity of incorporating more diverse training data and improving feature representation.
Q5. How sensitive the model performance is for different LTR model settings?
We tried Sigmoid, Tanh activation functions in the TMLP models to observe whether a non-linear function will impact the final result. Despite a little improvement, there is nothing statistically significant. We expect the small size of our test dataset and the big gap between training and testing data to make the results less sensitive to the structural perturbations on TMLP. The LTR models we trained are far from optimal, and better training data could improve the performance.
@@ 5.3. Discussion and Future Work
Our study represents a novel approach for search reranking. However, it is still preliminary and comes with several limitations. Therefore, it is necessary to examine those limitations and highlight our future directions.
@@
One limitation is that it has been mainly covered on the Android platform, and other platforms, such as iOS, desktop and web 999We’re now attempting to build a cross-platform execution proxy, following the design mentioned in the section 3.2.3. The key idea is to leverage platform-native API to capture accurate control hierarchies and expose UI-level interactions so that the technical concept in this paper can be naturally extended., should be studied in the future, considering their potential difference with Android, especially Accessibility API. Although we test diverse mobile application domains, the coverage is not exhaustive. Our study also shows that the exact performance (P@1, etc.) is domain-specific. This may affect the generalization of our findings, and the reader should consider this when interpreting the results. When applying similar techniques to other platforms or domains, the exact number (P@1 etc.) would differ.
@@
Although generative LLMs offer a quick start, the hallucination issue has heavily hindered the instruction extraction model, as noted by Ji et al. (Ji et al., 2023). However, during further analysis, we have found that such issue heavily relies on how we represent HTML data to LLMs and how data verification has been organized. In our recent experiment, we have improved the algorithm for extracting cleaned HTML and including it in the prompt with a clear XML tag, making LLMs recognize it as a code snippet instead of unstructured text. Together with the long context support of gpt4-turbo (GPT-4turbo, 2024), gpt4-o (GPT4o, 2024), a majority of instructions can be extracted successfully in the generative phase. In the grounding phase, we have reorganized the context in which the extracted instructions have been verified by the gpt4 models. We split the grounding phase into two subphases: the first phase focuses on the relevance between instructions and the given goal, and the next phase reexamizes whether the visual representation of instructions in HTML page is aligned with the goal. Through these refinements, we achieved a better result on our current dataset, which can serve as a more promising baseline in our future work. Looking forward, we would like to replace those two-phase solutions with a single LLM-extraction model: to be specific, we want to integrate generative LMs with multimodal document models such as XDoc (Chen et al., 2022), MarkupLM (Li et al., 2022b), LayoutLMv3 (Huang et al., 2022b), and LayoutLLM (Luo et al., 2024), aiming to enhance in-model extracting and grounding efficacy.
@@
A more tightly coupled integration of the UI context and the Action Prediction Model can further enhance the execution performance. Moreover, our current prompt-based method, which used the actionable regions as input, does not effectively represent semantic groups on the UI layer. For example, an input box and its accompanying label, intended for entering a username, should be semantically grouped as a single region. Another follow-up is to enhance the previous work (Hu, 2020), (Liu et al., 2018) to improve the UI contextual information passed to the Action Prediction Model. In our recent work, by using Dino v1 (Zhang et al., 2022) to understand UI segmentation and injecting into the UI contextual information, the execution module performs slightly better as knowing UI semantic groups. However, in the long term, a more promising direction is to follow the line of Ferret v2 (Zhang et al., 2024) by embracing the fine-grained UI representation learnt by Dino v2 (Oquab et al., 2023) into the vision part of the vision language model.
@@
Our current reranking model relies on simple feature engineering and does not use real-world training data. At the next step, we expect to acquire more real-world training data through strategic sampling. Rather than relying on manually engineered features, we could represent verification information, particularly the sequence of action screens and instructions, using more SOTA representation learning techniques. Using more realistic data and more advanced representation learning, we anticipate a substantial increase in the effectiveness of our proposed approach.
@@
Although verification agents help users avoid tedious manual verification, they might accidentally take harmful actions due to some misleading online content. In practice, we need to provide safeguards to prevent these risky actions from being taken by the execution agent. Ideally, a client environment that doesn’t impact the user’s real environment should be used. If not, we might allow each user to review and approve extracted instructions or at-risk actions before the automated execution. Another approach is to run server-side verification offline and display metadata about verification in the search results (e.g., ”verified for app1 v3 on android v11”), which allows personalized search results without realtime verification on the client side.
@@ 6. Conclusion
We propose reranking top retrieved results for ”How-to” search queries by promoting candidate web pages with verified executable instructions, which adds an additional layer of usefulness to a traditional search engine workflow. It ranks search results based on the actual completeness of candidate instructions, besides traditional metrics such as textual relevance and authority scores. The experimental results demonstrate our approach could further improve a very strong baseline search engine(i.e. Google).
This paper is a pioneer work in improving the reliability and usefulness of search results of online help resources for ”How-to” queries about software tasks. Along this direction, several promising future research and applications are on the horizon, like adapting the proposed workflow in figure 1 to support verification of other platforms (web, macOS, Windows, iPhone, etc.), or collecting user feedback to improve the reranking models and also improving the software agent’s ability to interpret and execute a wider range of instructions.
@@ Acknowledgements.
For the ”How-to” WeWeb dataset, we would like to appreciate our research collaborators: Sijia Zhong, Valentina Tang, Zoey Zhou, Jimmy Chen, Hanwen, Yang, and Chole Wong. Without their efforts, it would have been impossible to finish data collection on time. They also provided valuable suggestions and feedback to MagicWand. Additionally, we would like to thank Yaxuan Wang for her effort to refine the figures used to illustrate key ideas concisely and elegantly. Finally, we would like to thank our early contributors: Annabelle Miin, Anastasia Miin, Krish Pai, Rithvik Chavali, and Saket Pathak. They joined our research at the very beginning and worked tirelessly on data collection and verification. Their efforts encompassed sampling search queries, selecting relevant Android apps, and ensuring data accuracy through a rigorous two-stage verification process.References
@@
- GPT-4turbo (2024) GPT-4turbo. 2024. GPT-4turbo. OpenAI. https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo
- GPT4o (2024) GPT4o. 2024. GPT4o. OpenAI. https://platform.openai.com/docs/models/gpt-4o
@@
- Luo et al. (2024) Chuwei Luo, Yufan Shen, Zhaoqing Zhu, Qi Zheng, Zhi Yu, and Cong Yao. 2024. LayoutLLM: Layout Instruction Tuning with Large Language Models for Document Understanding. arXiv preprint arXiv:2404.05225 (2024).
@@
- Oquab et al. (2023) Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et al. 2023. Dinov2: Learning robust visual features without supervision. arXiv preprint arXiv:2304.07193 (2023).
@@
- Sarkhel et al. (2023) Ritesh Sarkhel, Binxuan Huang, Colin Lockard, and Prashant Shiralkar. 2023. Self-training for label-efficient information extraction from semi-structured web-pages. Proceedings of the VLDB Endowment 16, 11 (2023), 3098–3110.
- Sood and Dandapat (2023) Ojasvin Sood and Sandipan Dandapat. 2023. Problematic Webpage Identification: A Trilogy of Hatespeech, Search Engines and GPT. In The 7th Workshop on Online Abuse and Harms (WOAH). 126–137.
@@
- Zhang et al. (2022) Hao Zhang, Feng Li, Shilong Liu, Lei Zhang, Hang Su, Jun Zhu, Lionel M Ni, and Heung-Yeung Shum. 2022. Dino: Detr with improved denoising anchor boxes for end-to-end object detection. arXiv preprint arXiv:2203.03605 (2022).
- Zhang et al. (2024) Haotian Zhang, Haoxuan You, Philipp Dufter, Bowen Zhang, Chen Chen, Hong-You Chen, Tsu-Jui Fu, William Yang Wang, Shih-Fu Chang, Zhe Gan, et al. 2024. Ferret-v2: An Improved Baseline for Referring and Grounding with Large Language Models. arXiv preprint arXiv:2404.07973 (2024).